OpenStudy (anonymous):

Angles A and B are complementary angles in a right triangle. The value of csc A is 13/5. What is the value of cot A?

OpenStudy (anonymous):



OpenStudy (anonymous):

then tan A = 5/12 cot A = 12/5

OpenStudy (anonymous):

How did you get 12?

OpenStudy (anonymous):

\[12=\sqrt{13^{2}-5^{2}}\]

OpenStudy (anonymous):

since the two angles a and b are complementary then the third angle must be 90 degrees

OpenStudy (anonymous):

Oh, I see! thanks, that really clears it up.

OpenStudy (anonymous):

good luck :)